### v1

# MultiOCR-QA: Dataset for Evaluating Robustness of LLMs in Question Answering on Multilingual OCR Texts

Abstract.
Optical Character Recognition (OCR) plays a crucial role in digitizing historical and multilingual documents, yet OCR errors -imperfect extraction of the text, including character insertion, deletion and permutation- can significantly impact downstream tasks like question-answering (QA). In this work, we introduce a multilingual QA dataset MultiOCR-QA, designed to analyze the effects of OCR noise on QA systems’ performance. The MultiOCR-QA dataset comprises 60K question-answer pairs covering three languages, English, French, and German. The dataset is curated from OCR-ed old documents, allowing for the evaluation of OCR-induced challenges on question answering. We evaluate MultiOCR-QA on various levels and types of OCR errors to access the robustness of LLMs in handling real-world digitization errors. Our findings show that QA systems are highly prone to OCR induced errors and exhibit performance degradation on noisy OCR text.
Multilingual QA, OCR Text, Large Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; July 13–18, 2025; Padova, ITisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Question answeringccs: Information systems Content analysis and feature selection
1. Introduction
Optical Character Recognition (OCR) technology has played a crucial role in preserving and providing access to historical texts. Over the past decade, significant advancements in OCR have improved text recognition accuracy, leading to the development of large-scale digital libraries of historical texts (TERRAS, 2011). These libraries serve as valuable resources for researchers, historians, and the general public, granting access to old manuscripts, newspapers, and other archival materials. Historical documents hold a wealth of knowledge, offering insights into past events, cultures, and people. Many professionals such as historians, journalists, or sociologists rely on these collections for various research and analysis tasks.
Historical corpora have been widely used for numerous natural language processing (NLP) tasks, including Named Entity Recognition (NER), topic modeling, text classification, neural ranking, and understanding semantic changes in language over time (Ehrmann et al., 2020, 2023; Yang et al., 2011; Liebeskind and Liebeskind, 2020; Subramani et al., 2020). Among these, question answering (QA) provides a direct approach to extracting precise and relevant information from historical texts. Unlike broader text-processing tasks, QA focuses on retrieving specific answers to user queries, making it an essential tool for efficiently accessing historical knowledge.
Figure 1. An example of CorrectedOCR and RawOCR text from the MultiOCR dataset for the English language, highlighting different types of errors along with questions corresponding to this text. WER and CER denote Word Error Rate and Character Error Rate, respectively, indicating the level of errors in the text. The green highlights represent insertion errors, where IEO denotes Insertion Edit Operations - the number of insertions needed to transform RawOCR into CorrectedOCR. Similarly, red and blue highlights indicate deletion and substitution errors, with DEO and SEO representing Deletion Edit Operations and Substitution Edit Operations, respectively. The black boxes with numbers in the CorrectedOCR and RawOCR text correspond to the answers for each question in the paragraph.
Despite the advancements in OCR technology, several challenges persist. OCR-generated text, referred as RawOCR text in this paper, often contains errors due to the degraded conditions of historical documents. Factors like faded ink and paper, irregular fonts, physical damage, and printing inconsistencies cause recognition errors that negatively impact downstream NLP tasks such as information retrieval, machine translation, and QA systems. Since QA models heavily depend on the quality of the input text, errors in raw OCR text can significantly affect the accuracy and reliability of generated answers.
For instance, in German, a passage with OCR error ”Der Bericht der Lagsatzungsgesandtschaft wird verlesen undvon Hrn. Bürgermeistet Mousson als erstem Gesandten desStandes Zürich mit einigen Bemerkungen begleitet. [^1] English Translation: The report of the legislative mission is read out and accompanied by a few comments from Mr. Mayor Mousson, the first envoy of the Zurich state.” contains multiple insertion, deletion, and substitution errors, such as ”Lagsatzungsgesandtschaft” (should be ”Tagsatzungsgesandtschaft”) and ”Bürgermeistet” (should be ”Bürgermeister”). When a QA model encounters such a noisy OCR text, it may generate an incorrect answer. For example, given the question: ”Wer hat den Bericht der Tagsatzungsgesandtschaft verlesen?” (Who read the report of the parliamentary delegation?) the model incorrectly responds ”Der Bericht der Tagsatzungsgesandtschaft wurde von Hrn. Bürgermeistet Mousson verlesen.” (The report of the parliamentary delegation was read out by Mr. Mayor Mousson.) which retains the OCR error and potentially misleads the QA system. This example highlights how even minor OCR errors can significantly affect QA accuracy, leading to misleading or incorrect answers.
Although extensive research has focused on improving OCR accuracy and post-processing correction techniques, the specific impact of OCR noise on QA tasks remains underexplored. Previous studies have examined challenges related to OCR in information retrieval (IR) (de Oliveira et al., 2023), historical text processing (Piotrowski, 2012), and entity recognition (Grover et al., 2008). However, a systematic investigation of how OCR errors affect QA model performance is still lacking. Additionally, although large language models (LLMs) have been increasingly used to process OCR text, (Thomas et al., 2024; Madarász et al., 2024), their robustness in handling QA tasks with noisy OCR text remains an open research question.
In this paper, we address the critical gap in understanding how large language models (LLMs) perform in question-answering (QA) tasks when dealing with noisy OCR-generated text. Specifically, we analyze the robustness of LLMs in QA tasks using RawOCR text and introduce MultiOCR-QA  [^2] The dataset is freely available at: https://github.com/DataScienceUIBK/MultiOCR-QA, a new multilingual QA dataset constructed from the ICDAR dataset, covering English, French, and German. This dataset includes both RawOCR text (OCR-generated text with errors) and CorrectedOCR text (ground truth), allowing a direct comparison of QA performance under different text quality conditions. To generate contextually relevant question-answer pairs from historical text excerpts, we leverage instruction-fine-tuned LLMs. We then systematically evaluate the impact of different types of OCR errors—insertions, deletions, and substitutions—on QA model performance, offering new insights into the strengths and limitations of LLMs when processing noisy historical data. Figure 1 illustrates an example from the MultiOCR-QA dataset, showcasing various OCR errors and their impact on text accuracy. It presents QA pairs and responses using both CorrectedOCR and RawOCR text as context, demonstrating the effects of OCR noise on QA performance.
In summary, we make the following contributions in this work:
  • We introduce MultiOCR-QA, a new multilingual QA dataset from historical texts in English, French, and German, featuring both raw and corrected OCR text, allowing direct comparison under varying text quality conditions.
  • Using MultiOCR-QA, we then conduct a comprehensive evaluation of LLM robustness against noisy OCR text analyzing their impact on QA performance.
  • We also categorize different types and levels of OCR errors, evaluating their individual impact on QA performance, providing insights into how LLMs handle OCR-related challenges for different types of error.
Figure 2. The Pipeline for MultiOCR-QA generation: Arrows represent the output quantity based on the number of documents and callouts illustrate the statistics for English, French and German Language. Ps and QAs denote the number of paragraphs and question-answer pairs, respectively. Note, to supplement the data for English Language 10,000 questions from ChroniclingAmericaQA were included.
2. Related Work
Optical Character Recognition (OCR) technology has significantly contributed to the digitization of historical documents, enabling large-scale access to printed and handwritten archives. However, the accuracy of OCR output is highly dependent on factors such as document quality, font variations, language complexity, and historical text degradation. Early studies on OCR error analysis focused primarily on measuring character- and word-level error rates to assess their impact on text-based applications (Piotrowski, 2012; Hill and Hengchen, 2019).
Researchers have extensively studied the limitations of OCR for historical documents and its impact on information retrieval (IR). Croft et al. (1994) and Traub et al. (2015) examined how OCR errors reduce retrieval effectiveness. Chiron et al. (2017) found that 7% of the relevant documents were missed due to OCR misrecognition, demonstrating the risk of failure in matching noisy texts to user queries. While these studies highlight OCR challenges, they focus primarily on document retrieval and not on question answering (QA), which requires a more fine-grained understanding of text.
Beyond IR, OCR errors have been studied in multiple tasks, including named entity recognition (NER) (Hamdi et al., 2020, 2023), entity linking (Linhares Pontes et al., 2019), text classification (Zu et al., 2004; Vitman et al., 2022), topic modeling (Mutuvi et al., 2018; Zosa et al., 2021), document summarization (Jing et al., 2003), machine translation (Farooq1 et al., 2005; Khayrallah and Koehn, 2018), and document ranking (Giamphy et al., 2023). OCR noise has been shown to significantly degrade performance across these tasks. For instance, van Strien et al. (2020) demonstrated that low-quality documents negatively impact multiple tasks, including dependency parsing and sentence segmentation. Hamdi et al. (2022) found that 80. 75% of the named entities were misrecognized due to OCR errors, causing substantial drops in accuracy. Similarly, Hamdi et al. (2023) reported that the F1-score for NER drops from 90% to 50% when the character error rate increases from 2% to 30%. In topic modeling, Mutuvi et al. (2018) showed that OCR noise distorts the identification of key topics. For document retrieval, de Oliveira et al. (2023) analyzed performance degradation at different OCR error rates, noting that retrieval effectiveness begins to decline at a word error rate of 5% and worsens as the error rate increases. Giamphy et al. (2023) further examined the impact of different types of OCR noise on document ranking and advocated for developing more robust ranking methodologies.
Despite these insights into OCR’s effects on IR and NLP tasks, research on its impact on question answering remains limited. In the context of historical document collections, the only existing QA dataset, ChroniclingAmericaQA (Piryani et al., 2024), focuses primarily on creating a QA dataset from historical newspapers rather than systematically analyzing how different types of OCR errors affect QA performance. While studies on document retrieval and IR highlight OCR-related challenges, a comprehensive investigation into QA performance under different types and severity levels of OCR errors is still lacking. Our work fills this gap by introducing a multilingual QA dataset (MultiOCR-QA) and providing a detailed evaluation of large language models (LLMs) on the raw OCR text of MultiOCR-QA.
3. Methodology
To systematically investigate the impact of OCR errors on QA systems, we constructed MultiOCR-QA, a new multilingual QA dataset derived from historical texts processed with OCR. This section details the two main stages of the dataset creation pipeline: Data Collection and Question-Answer Generation. Figure 2 provides an overview of this process.
3.1. Data Collection
In this section, we describe the process of collecting documents to generate question-and-answer pairs to carry out our study. We utilized the ICDAR 2019 POST-OCR Text Correction dataset [^3] https://sites.google.com/view/icdar2019-postcorrectionocr (Rigaud et al., 2019) as our primary historical text source. We selected this dataset as it contains CorrectedOCR text (Ground Truth Text) along with RawOCR text. The dataset contains 22 million OCR-processed characters and corresponding ground truth text for ten European languages (English, French German, Finish, Spanish, Dutch, Czech, Bulgarian, Slovak, and Polish). In our study, we focused on English, French, and German.
Language Specific Data Collection: We extracted text in these three languages from the ICDAR 2019 dataset. The texts originally came from various historical document repositories.
  • English: The documents for the English language in the ICDAR 2019 dataset are sourced from IMPACT - British Library, comprising a total of 150 files.
  • French: For the French language, the ICDAR 2019 dataset provides a collection from three sources: the HIMANIS [^4] https://www.himanis.org Project, IMPACT - National Library of France, and the RECEIPT [^5] http://findit.univ-lr.fr/ dataset. The original dataset contained 2,800 files.
  • German: The German-language dataset includes the OCR-processed text from multiple sources, such as, front pages of the Swiss newspaper NZZ [^6] https://zenodo.org/records/3333627, IMPACT - German National Library, GT4Hist-dta19 dataset, GT4Hist - EarlyModernLatin, GT4Hist - Kallimachos, GT4Hist - RefCorpus-ENHG-Incunabula, and GT4Hist - RIDGES-Fraktur [^7] https://zenodo.org/records/1344132 (Springmann et al., 2018). The German dataset in ICDAR 2019 originally contained 10,032 files.
Language Verification and Filtering: Prior to QA pair generation, we preprocessed the ICDAR dataset to ensure that each document contains text in the correct target language. We applied langdetect [^8] https://github.com/Mimino666/langdetect library to detect the language of documents. The analysis revealed that some documents labeled as English, French and German were actually in other languages, particularly Latin. To maintain dataset integrity, we removed non-target language documents, resulting in the following reductions: We removed non-English documents, reducing the number of documents in the English dataset to 141. Similarly, we discarded non-French documents, reducing the dataset to 1,713 French-language files, and eliminating 1,086 Latin-language files. Finally, for German, we removed Latin or other non-German files and retained a total of 9,075 German-language files.
Furthermore, the ICDAR dataset, originally intended for post-OCR correction, contained special alignment symbols (e.g., @, #) to map the RawOCR text to its ground-truth counterpart. We removed the alignment symbols from the ground-truth text before generating questions. We also excluded files where the ground-truth text was missing, resulting in the removal of 16 files for English, 3 files for French, and none for German. This preprocessing step ensured that all QA pairs were generated from text that had both CorrectedOCR text and RawOCR text [^9] We will sometimes use CorrectedOCR and RawOCR to denote the text that was subject to post-OCR error correction and one without such correction, respectively..
3.2. Question-Answer Generation
To construct the multilingual QA dataset, we opted for automatic QA pair generation, as manual dataset creation would require substantial human resources. To achieve this, we instruction fine-tuned a pretrained LLM for each target language to generate QA pairs from CorrectedOCR text.
While LLMs are pretrained on diverse NLP tasks, they typically generate a variety of question types, including non-factoid and open-ended questions. Since our goal is to develop a factoid QA dataset, we fine-tuned the models using language-specific QA datasets to ensure the generation of structured, precise, and factual question-answer pairs.
Instruction fine-tuning enhances both the capabilities and controllability of LLMs (Zhang et al., 2023). Fine-tuning instruction-based datasets across multiple languages allows the model to generalize across different question-answering styles, ensuring that the generated questions remain relevant even when dealing with language-specific variations. We opted to finetune LLama-3.1-70B instruct model (Dubey et al., 2024) separately for each language using widely adopted QA datasets.
For fine-tuning the model in English, we used the SQuAD v1 dataset (Rajpurkar et al., 2016). We randomly selected 2,067 paragraphs and 10,570 questions from the development set and 3,000 paragraphs and 13,894 questions from the test set. In total, we fine-tuned the LLaMA-3.1-70B instruct model for English QA dataset generation using 5,067 paragraphs and 24,464 questions. Similarly, for the French language, we fine-tuned the model on the FQuAD dataset (d’Hoffschmidt et al., 2020), utilizing both the validation and training sets, comprising 5,689 paragraphs and 23,919 questions. For the German language, we fine-tuned the model on the GermanQuAD dataset (Möller et al., 2021) using the training and validation sets, which consisted of 3,014 paragraphs and 13,722 questions. This fine-tuning step ensured that the model accurately generated factoid-style QA pairs, reducing instances of open-ended or conversational questions.
After instruction fine-tuning, we used the fine-tuned model for each language to generate questions for the preprocessed dataset prepared in the initial step. To systematically generate high-quality QA pairs, we employed a two-step prompt-based approach: Answer Extraction and Question Generation for the Extracted Answers.
Answer Extraction: The model was first prompted to extract multiple candidate answer spans from a given passage. These spans included entities, numbers, dates, locations, and key phrases that could serve as factual answers. The following prompt was used for the extraction of the candidate answer.
English Answer Extraction Prompt System: You are an expert at extracting key information from text. Your goal is to identify spans of text that are likely to serve as answers to potential questions based on the input passage. Focus on meaningful, distinct, and diverse snippets such as entities, nouns, verbs, adjectives, numbers, dates, and phrases. Avoid redundancy and ensure the answers are diverse, representing key information in the passage. User: Given the passage below, extract several candidate spans that are likely to be answers to potential questions. Write only the extracted answers, separated by a semicolon (;). Passage : {context}
After generating the candidate answers, we checked for duplicated answer spans. If the answers were duplicate, we removed them and retained only unique answers. This prompt was applied uniformly across all three languages, with translations adapted to each language.
Question Generation: Following answer extraction, the extracted answer spans were re-fed into the model, and it was prompted to generate questions that align with each answer while maintaining contextual relevance. The generated questions were structured to be standalone, well-formed, and factually grounded in the passage. The following Prompt was used for Question Generation from the Extracted Answers.
English Question Generation Prompt System: You are an expert at generating standalone questions based on the provided passage. Your goal is to create a clear, relevant, and self-contained question that aligns with the information in the passage. The question should not explicitly reference the passage or require additional information to be understandable. Ensure the question is concise, well-structured, and meaningful. User: Based on the passage below, please generate a question that is relevant to the information provided. The question should be standalone, clear, and understandable without referencing the passage directly. The answer to the question should be [answer]. Passage : {context}
Using this approach, we generated 941 questions for English, 10,522 questions for French, and 44,607 questions for German.
Dataset Filtering: After generating the dataset for each language, we applied additional filtering steps to ensure quality and consistency. Specifically, we removed questions that did not end with a question mark, duplicate questions, and questions with excessively long answers. Since LLM-generated datasets may contain hallucinated long answers, we applied this additional filtering by removing excessively long answers.
As a result, we removed 66 questions for English Language, 530 questions for French Language, and 7,210 questions for German Language. This final filtering step ensured that the MultiOCR-QA dataset consisted of concise, well-structured, and factually accurate question-answer pairs.
4. Dataset Analysis
After applying all the filtering steps, we obtained the final dataset, comprising 50,079 question-answer pairs. Due to the relatively low number of English-language documents in the ICDAR 2019 dataset, the initial MultiOCR-QA dataset contained only 875 QA pairs for English. To balance the distribution of QA pairs across languages, we incorporated 10,000 additional QA pairs from the ChroniclingAmericaQA dataset (Piryani et al., 2024). That dataset, sourced from the American historical newspaper collection called Chronicling America [^10] https://chroniclingamerica.loc.gov/about/, spans over 120 years from 1800-1920. It includes both ground-truth and RawOCR text, making it particularly relevant to our study. Consequently, the final MultiOCR-QA dataset consists of documents from historical newspapers, books, and centuries-old scripts, with 10,875 QA pairs in English, 10,004 in French, and 39,200 in German. The dataset statistics, including average paragraph length, question length, and answer length, are presented in Table 1.
Table 1. Basic statistics of the MultiOCR-QA dataset, including question-answer (QA) pair count, paragraph count, and average text lengths across languages.
English French German
#QA pairs 10,875 10,004 39,200
#Paragraphs 6,525 1,670 9,075
Average CorrecteddOCR paragraph length (words) 219.09 297.53 212.86
Average RawOCR paragraph length (words) 233.25 335.73 193.23
Average question length (words) 10.98 8.73 8.08
Average answer length (words) 2.05 3.12 5.63
Average questions per paragraph 1.67 5.99 4.32
Figure 3. Statistics of insertion, deletion, and substitution errors for each language, categorized into low, medium, and high noise levels.
4.1. Quantifying and Filtering OCR Noise
To assess the impact of OCR noise on QA quality, we quantified the noise level in the RawOCR text using two standard metrics: Character Error Rate (CER) and Word Error Rate (WER). CER measures the proportion of character-level errors in the raw OCR text compared to the ground truth text. It is computed as the number of insertions, deletions, and substitutions (including spaces) required to transform the RawOCR text into its correct form. WER quantifies word-level discrepancies, representing the proportion of words that require modifications (insertions, deletions, or substitutions) to match ground truth text. Both CER and WER were computed using the Levenshtein distance (Miller et al., 2009), which determines the minimum number of edits needed to correct the OCR-generated text. A high CER but low WER suggests that errors are concentrated within a few words (e.g., spelling variations), whereas a high WER indicates distortions across multiple words, significantly affecting readability.
Outlier Detection: To ensure a reliable analysis, we applied the Interquartile Range (IQR) method to detect and remove outliers in CER values. Outliers were defined as CER values below Q1 - 1.5 * IQR or above Q3 + 1.5 * IQR, where Q1 and Q3 represent the 33th and 66th percentiles of the CER distribution, respectively, and IQR is the difference between Q3 and Q1. This filtering resulted in the removal of 1,056 English, 351 French, and 2,423 German paragraphs.
Following outlier removal, we categorized the remaining paragraphs into three noise levels based on CER percentiles for each language. Documents with CER below the 33rd percentile were classified as ”low noise,” those between the 33rd and 66th percentiles as ”medium noise”, and those above the 66th percentile as ”high noise.” The specific CER thresholds for each category and language were as follows:
[…]
In addition to CER-based classification, we analyzed the distribution of three specific OCR error types: insertions, deletions, and substitutions. Each error type was categorized separately using a percentile-based approach, allowing for a more detailed examination of the nature and severity of OCR distortions. To further investigate OCR noise patterns, we classified insertion, deletion, and substitution errors into low, medium, and high noise levels. As illustrated in Figure 3, the distribution of these error types varies significantly across languages, reflecting differences in OCR quality and text processing challenges in English, French, and German. Additionally, Table 2 presents the statistical characteristics of the distribution of OCR error metrics across languages, including CER, WER, and edit operations, providing further insights into the OCR noise characteristics.
Table 2. OCR Error Statistics across Languages
Metric English French German
Character Error Rate (CER)
Mean 0.0988 0.0519 0.2816
Median 0.0826 0.0440 0.2592
Word Error Rate (WER)
Mean 0.2587 0.1904 0.8713
Median 0.2319 0.1760 0.8730
Edit Operations
Mean Substitutions 74.66 31.51 240.23
Median Substitutions 61.00 21.00 230.50
Median Deletions 11.74 41.09 85.09
Median Deletions 7.00 25.00 39.00
Median Insertions 39.02 20.90 82.60
Median Insertions 30.00 11.00 80.00
4.2. Human Evaluation
To assess the quality of the MultiOCR-QA dataset, we also conducted a manual evaluation study. A total of 100 question-answer pairs were randomly selected for evaluation from each language. We asked native speakers for each language, to rate the questions. The evaluation followed a five-point rating scale (1 to 5), where 1 indicated very poor and 5 indicated very good. The assessment was based on four key criteria: (1) Question Readability, assessing grammatical correctness and fluency; (2) Answer Readability, ensuring coherence and grammatical correctness; (3) Relevance, verifying whether the generated question aligns with the passage; and (4) Answer correctness, confirming the accuracy and completeness of the answer.
Each question was rated by an annotator, and the final scores for each criterion were averaged to obtain an overall evaluation. The results of the human evaluation are presented in Table 3. The evaluation showed high scores for both question and answer readability for the English and French language. The scores for relevance and answer correctness were moderate for these languages, indicating that most of the questions generated were contextually appropriate and that the answers provided were generally correct. But the results for German language were very low depicting that the CorrectedOCR text has a huge number of OCR errors so the errors are carried out further in the generated dataset. Such levels of errors are common when digitizing old historical texts due to the physical damage, irregular fonts, or faded ink of original document (Jatowt et al., 2019).
Table 3. Human evaluation results of MultiOCR-QA dataset.
Language Question Readability Answer Readability Relevance Answer Correctness
English 4.13 4.03 3.33 3.75
French 4.16 4.14 3.75 3.30
German 2.80 1.92 1.64 1.59
5. Experiments and Results
In this section, we conduct a comprehensive analysis of MultiOCR-QA from several perspectives. First, we evaluate the performance of MultiOCR-QA across various LLMs, comparing different model families and sizes to assess their effectiveness in handling OCR-generated text. Second, we investigate the impact of OCR errors on QA performance, focusing on three primary error types: insertion errors, deletion errors, and substitution errors. By analyzing these aspects, we aim to quantify the robustness of LLMs in processing noisy OCR text and provide insights into the challenges associated with question-answering on historical documents.
5.1. Experimental Settings
We conducted experiments using multiple large language models (LLMs), including Qwen2.5 7B (Yang et al., 2024), LLaMa 3.1-8B (Dubey et al., 2024), Gemma-2-27B (Team et al., 2024), Mixtral 8×7B (Jiang et al., 2024), LLaMA 3.3-70B (Dubey et al., 2024), and Qwen2.5 72B (Yang et al., 2024). These models span different architectures and parameter sizes, allowing for a comprehensive comparison on OCR text.
Traditionally, QA systems are evaluated using Exact Match (EM). However, these metrics can be insufficient for LLMs, as models often generate verbose responses, leading to low EM scores even when the correct answer is included in the response. To address this limitation, we evaluate MultiOCR-QA using BERTScore (Yuan et al., 2021) alongside EM. Additionally, we introduce another evaluation metric, Contains, to better assess MultiOCR-QA performance. Contains measures the extent to which the ground truth is present in the response generated by the model, regardless of the verbosity. We will apply these metrics to evaluate the QA results on RawOCR texts and then on CorrectedOCR texts used as context.
5.2. Experimental Results
Tables 4, 5, and 6 present the impact of OCR errors on the performance of various LLMs in question-answering tasks for English, French, and German texts.
Table 4 presents the performance of LLMs on English text using both CorrectedOCR (CP) and RawOCR (RP) paragraphs. Across all models, the transition from CP to RP negatively impacts performance. The best-performing model on CP is Gemma-2 27B with a BERTScore of 62.90, followed by LLaMA-3.3-70B at 60.71. When switching to RP, Gemma-2 still achieves the highest BERTScore 59.70, but with a 5.09% drop, highlighting its robustness. The lowest-performing model is Mixtral 8×22B, showing the most significant impact of OCR errors. Qwen-2.5 72B model achieves the best performance in CP for the Contains metric of 59.95, indicating its strong retrieval ability in clean text. However, it experiences a 9.48% drop in RP, suggesting moderate sensitivity to OCR errors. The lowest performing model in RP is again Mixtral 8×22B for Contains metric which sees an 11.26% decrease, indicating its greater vulnerability to noisy text. The Exact Match (EM) metric shows the steepest decline across models, emphasizing that OCR errors severely impact the models’ ability to generate precise answers.
Table 5 reports the performance of LLMs on French text, again comparing CorrectedOCR and RawOCR paragraphs. The same as in English, the models perform better with CorrectedOCR, confirming and quantifying the negative influence of OCR errors on the QA accuracy. Gemma-2 27B model achieves the highest BERTScore of 76.51 on CP, demonstrating its strong ability to capture semantic similarity. Despite the 1.46% drop, it still maintains the highest performance with a BERTScore of 75.39 on RP, indicating robustness to OCR noise. Mixtral 8×22B model shows the smallest drop of 1.05%, but its overall score remains lower than the ones for the other models. For Contains, Qwen-2.5 72B achieves the highest score on CP 57.55, highlighting superior retrieval performance on clean text. However, it experiences a 19.90% drop in RP, reinforcing its vulnerability to OCR errors. LLLaMA-3.1 8B also struggles, with a 19.63% decline, showing difficulty in retrieving spans from noisy text. In terms of EM, Gemma-2 27B outperforms all models with an EM of 17.46 on CP. Although it drops by 15.23%, it still maintains the best EM value 14.80 in RP. Mixtral 8×22B model is the most affected in EM, dropping by 27.99%, suggesting that OCR noise drastically reduces its accuracy in generating exact answers.
Table 4. Performance of LLMs on English Language: Comparison Using CorrectedOCR Paragraphs (CP) and RawOCR Paragraphs (RP) as Context. Red numbers indicate the percentage decrease in performance with RP. Bold values highlight the highest performance for each metric with CP, while underlined values denote the best performance for each metric with RP.
Model Parameter BERTScore Contains EM
Qwen-2.5 (CP) 7B 49.58 57.36 8.50
[…]
LlaMA-3.1 (CP) 8B 50.18 59.27 0.344
[…]
Gemma-2 (CP) 27B 62.90 53.16 19.05
[…]
Mixtral (CP) 8x22B 40.81 57.89 1.92
[…]
LlaMA-3.3 (CP) 70B 60.71 53.39 11.22
[…]
Qwen-2.5 (CP) 72B 49.64 59.95 11.53
[…]
Table 5. Performance of LLMs on French Language: Comparison Using CorrectedOCR Paragraphs (CP) and RawOCR Paragraphs (RP) as Context. Red numbers indicate the percentage decrease in performance with RP. Bold values highlight the highest performance for each metric with CP, while underlined values denote the best performance for each metric with RP.
Model Parameter BERTScore Contains EM
Qwen-2.5 (CP) 7B 73.12 54.82 11.38
[…]
LLaMA-3.1 (CP) 8B 70.16 51.35 0.64
[…]
Gemma-2 (CP) 27B 76.51 52.58 17.46
[…]
Mixtral (CP) 8x22B 68.65 48.32 0.30
[…]
LLaMA-3.3 (CP) 70B 72.50 54.00 4.41
[…]
Qwen-2.5 (CP) 72B 73.26 57.55 10.62
[…]
Table 6. Performance of LLMs on German Language: Comparison Using CorrectedOCR Paragraphs (CP) and RawOCR Paragraphs (RP) as Context. Red numbers indicate the percentage decrease in performance with RP. Bold values highlight the highest performance for each metric with CP, while underlined values denote the best performance for each metric with RP.
Model Parameter BERTScore Contains EM
Qwen-2.5 (CP) 7B 62.76 15.88 0.389
[…]
LLaMA-3.1 (CP) 8B 63.87 15.36 0.457
[…]
Gemma-2 (CP) 27B 67.07 11.56 2.691
[…]
Mixtral (CP) 8x22B 60.87 11.23 0.078
[…]
LLaMA-3.3 (CP) 70B 63.82 17.68 0.553
[…]
Qwen-2.5 (CP) 72B 63.25 16.43 0.699
[…]
Figure 4. BERTScore of different error types on Low, medium and High categories for each Language in MultiOCR-QA dataset.
Table 6 examines the performance of the model in the German text, revealing the most significant performance decline among the three languages. Unlike English and French, German exhibits the largest performance drop due to its lower OCR quality, and the models struggle more with its linguistic structure. Among the models evaluated, Gemma-2 27B achieves the highest BERTScore of 67.07 on CP, confirming its strong ability to capture semantic similarity in clean text. It also maintains the best performance on RP (63.78); however, it still experiences a 4.91% decrease, highlighting the adverse effects of OCR errors. In contrast, LLaMA-3.1 8B and LLaMA-3.3 70B show the biggest BERTScore drop (8.74% and 8.43% respectively), suggesting that these models struggle more with OCR noise. For the Contains metric, LLaMA-3.3 70B achieves the highest Contains score on CP, making it the most effective for retrieving relevant information in clean text. However, all models suffer a severe drop in Contains when moving to RP, with reductions exceeding 65%. Qwen-2.5 7B has the worst drop 66.40%, indicating that it faces challenges to retrieve information from noisy text. In terms of EM, Gemma-2 27B achieves the highest EM score on CP 2.69, while Mixtral 8×22B has the lowest EM 0.078. EM scores drop drastically across all models, the largest decrease is 79.04% for Gemma-2, reinforcing that word-level distortions from OCR errors make exact answer matching nearly impossible. The Mixtral model drops 72.74% in EM, making it highly unreliable for exact answers in noisy OCR text.
Summary of findings: OCR errors consistently degrade the performance of the models in English, French and German texts, resulting in maximum drop of 5.09%, 1.75%, and 8.74% in BERTScore, respectively. The most severe impact was observed in German due to the lower quality of the OCR and the complex linguistic structure. While larger models like Gemma-2 27B and LLaMA-3.3 70B demonstrate greater resilience, all models suffer substantial declines in Contains and Exact Match (EM) metrics, highlighting their weakness in retrieving and generating precise answers from noisy text. Gemma-2 27B consistently outperforms others, maintaining the highest BERTScore and EM across all languages, but still experiences notable degradation in noisy conditions. Mixtral 8×22B emerges as the most vulnerable, exhibiting the lowest absolute performance and struggling particularly with exact answer generation.
5.3. Performance based on Different Error Types
In this section, we conduct an in-depth analysis of the impact of different types of OCR errors. Insertion, deletion, and substitution on QA systems. We use Gemma-2 (27B) for this analysis, as it was found to consistently outperform the other models across English, French, and German. As detailed in Section 4.1, each error type is categorized into three levels: Low, Medium, and High, where Low represents minimal presence and High indicates the most frequent occurrence of a particular type of error. We evaluated MultiOCR-QA’s performance across these categories, as illustrated in Figure 4.
Insertion errors introduce extraneous characters or words, leading to moderate performance degradation. At low and medium insertion levels, the effect on BERTScore remains relatively minor, suggesting that small insertions do not always disrupt semantic meaning. However, at high insertion levels, performance drops sharply, indicating that excessive insertions impair both readability and semantic coherence.
Deletion errors impact sentence coherence and factual consistency, especially when they corrupt or remove key words or essential contextual phrases. Although the impact is less pronounced at lower levels, it escalates sharply as the frequency of deletions increases. At higher levels of deletion error, the degradation in BERTScore is similar to that seen with substitution errors, highlighting how missing characters or words disrupt structured text.
Substitution errors exhibit the most severe impact on QA performance in English and French, causing the steepest decline in BERTScore as their frequency increases. Since these errors modify characters within words, they often alter word meaning and disrupt sentence structure, making them highly detrimental to text comprehension.
However, in German, substitution errors appear to be less disruptive than in English and French. This can be attributed to the compound word structure in German, where minor substitutions can still preserve some semantic similarity. In contrast, deletions or insertions tend to fragment meaningful lexical units, making them more impactful in German than in other languages.
Summary of findings: Across languages, the results reveal that English and French exhibit similar degradation patterns, with BERTScore progressively decreasing as the OCR error frequency increases. However, in German, the sharpest decline is observed across all error types, particularly for substitutions and deletions. This suggests that OCR noise in German is more detrimental, probably because the older scripts have content in old German where characters such as (long s) are used instead of ”s”, which can be often misread as ”f” or ”l”.
The results indicate that substitution errors are the most disruptive in English and French, while German is more affected by deletions due to its compound word structure. Insertion errors generally cause moderate degradation, but severe performance drops occur at high error levels. Overall, German experiences the highest performance drop, reinforcing its greater vulnerability to OCR distortions and highlighting the need for effective OCR correction strategies.
5.4. Performance based on Pre-Processed and Post-Processed OCR Text
In this section, we conduct an additional study to assess the impact of OCR on QA through pre-processing and post-processing Raw OCR text. In this experiment, we focus exclusively on paragraphs and their corresponding questions derived solely from the ICDAR dataset for the English language, comprising 83 paragraphs and 513 questions. We chose this subset of paragraphs and questions because its OCR ground truth is manually curated, offering a robust basis for comparison, unlike the CorrectedOCR text in the ChroniclingAmerica QA, which has been refined using GPT 3.5 (Brown et al., 2020). Our approach involves two strategies: pre-processing, where RawOCR text is corrected using the Gemma-2 27B model prior to answering the questions, and post-processing, where answers generated from RawOCR text are subsequently corrected. Gemma-2 27B was used in this experiment as it is identified as the best performing model among all other LLMs as demonstrated in Section 5.2. The results for each strategy are presented in Table 7.
Table 7. Performance Metrics of QA Systems Using Pre-Processed and Post-Processed OCR Text. Red indicates the percentage drop in performance when using different variants of paragraph used as context, compared to Corrected Paragraph (Ground Truth).
Approach BERTScore Contains EM
Corrected Paragraph (Ground Truth) 63.99 58.09 16.76
[…]
The results in Table 7 indicate that using ground truth paragraph as context gives better performance metrics, with a BERTScore of 63.99, Contains of 58.09, and EM of 16.76, compared to using RawOCR text or LLM corrected passage as context. The comparative analysis reveals that pre-processing the RawOCR text significantly enhances QA performance, as evidenced by higher BERTScore, Contains, and EM scores. This suggests that correcting OCR errors prior to QA processing preserves the semantic integrity of the text, thereby facilitating more accurate answer generation. However, it shows a percentage drop of 7. 28% compared to the Ground Truth BERTScore. In contrast, post-processing the generated answer is less effective, resulting in the highest drop of 12.33% and the lowest BertScore among others. This is likely because that approach does not include any contextual reference and leads to generation of corrected words that have limited relation to context.
These findings emphasize the importance of integrating OCR error correction early in the QA pipeline to improve the reliability of QA systems, especially when dealing with historical texts or other archival materials. However, given the huge collections of digitized content with vastly varying levels of OCR quality that the current memory institutions (archives, libraries, museums, etc.) held, the correction cost and effort would be enormous. It is also difficult to correct the queried texts at inference time as this would also introduce computational cost and latency in online systems. Therefore, more robust QA systems that are aware of OCR errors and capable of predicting correct answers based on contextual information are required.
6. Conclusion
In this paper, we introduce MultiOCR-QA, a multilingual question-answering dataset for historical texts processed by OCR. Our dataset is unique as it is specifically designed to assess the impact of OCR errors on QA system performance, enabling a systematic analysis of how different error types—insertions, deletions, and substitutions- affect the accuracy and robustness of large language models. By incorporating both CorrectedOCR (clean text) and RawOCR (OCR-ed text), MultiOCR-QA facilitates comparative assessments that highlight the challenges posed by OCR noise in historical document processing. Our evaluation of multiple state-of-the-art LLMs demonstrates that OCR errors significantly degrade QA performance, with the most pronounced effects observed in languages with higher OCR error rates. Although larger models such as Gemma-2 27B and Qwen-2.5 72B exhibit greater resilience to OCR noise, smaller models and low-resource QA systems show a more substantial performance drop. These findings underscore the need for OCR-aware training strategies, particularly for historical document processing and multilingual QA tasks.
Use cases: The MultiOCR-QA dataset offers a unique resource to advance research on OCR-aware QA and studying QA on noisy OCR text, making it useful in several ways; It can be used to train LLMs to improve error correction capabilities and enhancing robustness against OCR inaccuracies while preserving the archaic language structure. It can be also used to expand LLMs’ multilingual processing abilities by training on multilingual OCR texts, improving accuracy in low-resourced languages.
Limitations and Future work: While MultiOCR-QA includes English, French and German, it does not encompass low resource languages or scripts such Latin, Finish and others. Future research should incorporate low-resourced languages to improve generalization and greater applicability across diverse languages. Additionally, methodologies that not only remove OCR errors but also preserve the original structure of documents could be applied.